# Mid-Infrared Dust Evolution and Late-time Circumstellar Medium Interaction in SN 2017eaw

### v1

@@ 3.5 Dust Modeling Results
The Cycle 1 JWST observations were previously modeled for the optically thin and dusty sphere case in M. Shahbandeh et al. (2023). M. Shahbandeh et al. (2023) report only a total dust mass rather than separate mass components as we do here. However, we are able to reproduce all the dust properties for the dusty sphere of silicate dust case reported in M. Shahbandeh et al. (2023) by assuming since . For the optically thin case, we are unable to replicate the temperature of the hotter dust component as M. Shahbandeh et al. (2023), though we reproduce the mass of the dust (again assuming ). However, we find that our Cycle 1 value is consistent within error with our Cycle 2 value, which is well constrained by the NIR data.
@@ 3.6 Dust Geometry
First, to determine the robustness of any modeled dust radii measurements, we apply the same check as S. Zsíros et al. (2024) and calculate the optical depth as follows (L. B. Lucy et al., 1989):
@@ 3.7 Spectral Evolution
Recent work by L. Dessart et al. (2025) presents models of a SN II at 1000 days post-explosion that is CSM-interacting and also contains a small mass of dust in the cold dense shell. They find there is no significant dust attenuation effect for dust interior to the ejecta due to the small angle of the inner dust relative to the emitting region of the outer ejecta. Further, in contrast to non-interacting SNe, in the L. Dessart et al. (2025) CSM-interacting SN model interior dust has no impact on the line strengths, or the hydrogen-oxygen line ratios, due to the fact that 99.7% of the model emission is from the outer ejecta. Indeed, K. E. Weil et al. (2020) note no significant signs of dust attenuation in SN 2017eaw at 900 days post-explosion. At 2668 days post-explosion, the velocity of the ejecta (7000 km/s) is similar to that of the 1000 day model (8000 km/s) in L. Dessart et al. (2025) and it is possible that the observed line attenuation is not due to interior dust. Although this physical picture may still be valid, we caution against direct comparisons given that both the inner and outer regions have evolved for an additional 1700 days. The velocity of the ejecta at 2668 days suggests that the outer regions of the ejecta have somewhat slowed due to CSM interaction and the conclusions from 1000 days may no longer be valid. L. Dessart et al. (2025) also investigate the effect of , , and of dust in the ejecta at a velocity coordinate of 8000 km/s. These models produce H profiles very similar to the one at 2668 days. The model profiles include a dip in the center (near rest wavelengths) due to the increased optical depth of the limbs of the shell. We refrain from definitively linking a similar feature in the observed H profile with dust given the low SNR of the spectrum. Further modeling of the effects of dust in CSM-interacting SNe II at late times (1000 days post-explosion) is needed to understand the evolution of H and the strongest oxygen lines.
@@ 4.1 Collisionally Heated Pre-existing Dust
The mid-infrared SED of SN 2017eaw is somewhat similar in shape to that of SNe 1987A and 1980K at significantly later phases, see Figure 9. In the case of both SNe 1987A and 1980K, the 8-20 m emission is consistent with 150-180K silicate dust, remarkably similar to SN 2017eaw. Notably the mid-infrared flux in SN 2017eaw, which is primarily dominated by the cool dust component, is almost identical in shape to Spitzer IRS spectra of SN 1987A between 6000 and 8000 days post-explosion. This component in SN 1987A has been linked to the collisional heating of the equatorial ring (E. Dwek et al., 2010; R. G. Arendt et al., 2016, 2020), and a similar scenario was suggested for SN 1980K (S. Zsíros et al., 2024). The mass of dust in the 160K component of SN 2017eaw is 1 order of magnitude larger than observed in SN 1987A and 1 order of magnitude smaller than observed in SN 1980K.
To determine if it is plausible for the cooler component of SN 2017eaw’s mid-infrared dust to be collisionally heated via interaction between the ejecta and CSM, we follow the method of O. D. Fox et al. (2010) for estimating the mass of dust processed by the forward shock (see also O. D. Fox et al., 2011; S. Tinyanont et al., 2016; S. Zsíros et al., 2022, 2024). Any pre-existing dust must reside outside the evaporation radius, inside which the peak luminosity of the SN will have destroyed any pre-existing dust grains. Assuming the temperature and peak bolometric luminosity measured by T. Szalai et al. (2019b), 14,000K and 1043 erg/s (rounded from L erg/s) respectively, the evaporation radius (Revap) is R⊙. Revap is significantly less than the ejecta radius at 2,000 days. In both JWST epochs, the SN ejecta has far surpassed the evaporation radius and could feasibly be interacting with CSM containing pre-existing dust.
@@
where is the shock velocity, is the time since explosion, and is the dust grain size. Similar to S. Zsíros et al. (2024), we use 5,000 km s-1 and 15,000 km s-1 and 0.005 and 0.1 m (we assume m in our dust modeling) as our lower and upper bounds, respectively. Assuming these values, the range of dust masses that could be collisionally heated is . The total dust masses observed in both Cycle 1 and 2 are in the middle of this range. We note that the velocity of the ejecta measured from the nebular spectra, 7000 km s-1, with a 0.1 m dust grain could result in a collisionally heated dust mass of . This result suggests that the majority of the dust observed in the mid-infrared could reasonably be collisionally heated pre-existing dust.
@@ 4.2 Radiatively Heated Pre-existing Dust
A simple IR echo model, assuming the light from the SN explosion excites pre-existing dust, places the echo radius at , where is the duration of the light echo (M. F. Bode & A. Evans, 1980; E. Dwek, 1983). Assuming a lower limit of days in order for the echo to still be detectable in the Cycle 2 mid-infrared observations, this gives , significantly above the outer dust radii given by the non-optically thin models in Section 3.4. When the ejecta is between the evaporation and the light echo radii, as is the case for SN 2017eaw, the luminosity from CSM-ejecta interaction heats the pre-existing grains creating a CSM echo (O. D. Fox et al., 2010, 2011)888Sometimes referred to as a circumstellar shock echo. M. Shahbandeh et al. (2023) found that for SN 2017eaw the optical luminosity necessary to heat the grains to the temperature of the cool dust component in the Cycle 1 data exceeds the observed optical luminosity. Unsurprisingly, we find this to be the case for the Cycle 2 dust as well.
@@ 4.4 Implications for Dust Formation in CCSNe
If the mid-infrared dust emission in SN 2017eaw is primarily due to pre-existing dust, then its location on the dust formation timeline may be significantly different than shown in Figure 11. It is possible that many of the early time mid-infrared dust measurements of CCSNe are similarly contaminated with pre-existing dust. Late time dust mass measurements of SN 1987A were done in the far-IR and sub-mm and probed dust significantly colder than can be observed with JWST (M. Matsuura et al., 2011; R. Indebetouw et al., 2014; M. Matsuura et al., 2015). For SN 1980K, reported dust measurements were measured by modeling the dust attenuation on optical spectral lines. The mid-infrared SED of SN 1980K reveals 100 times less dust than indicated by the line profiles (S. Zsíros et al., 2024). This suggests that the majority of the dust in SN 1980K is also too cold to be observed by JWST/MIRI. The same might be true for SN 2017eaw but the signal to noise of the recent spectra makes modeling of the line profiles difficult.
@@ 6 Acknowledgments
This work is based in part on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program GO3295 and GO2666.
@@
In order to measure the flux in the MIRI images, we attempted several different photometric techniques which we compared to the published Cycle 1 photometry in M. Shahbandeh et al. (2023). In this work, we report JWST MIRI PSF photometry of SN 2017eaw as this methodology resulted in values that are most consistent with previously published photometry. M. Shahbandeh et al. (2023) did PSF photometry on the stage 2 products, a method similar to the one we use for our photometry in Section 2.1. As shown in Figure 12 (purple diamonds and pink squares), the most significant offsets between our space_phot PSF photometry and the published photometry are at 18 and 21 µm, both of which are filters where the photometric calibration maps were significantly updated999This work uses version 0056, details can be found at https://jwst-crds.stsci.edu/browse/jwst_miri_photom_0056.rmap between the publication of M. Shahbandeh et al. (2023) and the completion of this work, therefore this offset is unsurprising.
@@
Due to the discrepancy between the aperture and PSF photometry regardless of methodology, and the existence of published JWST/MIRI PSF photometry of SN 2017eaw, we opt to report only the PSF photometry discussed in Section 2.1 in this work. Importantly, we do find that, regardless of the photometric method, the total mid-infrared flux of SN 2017eaw has decreased from Cycle 1 to Cycle 2. We caution that the aperture and PSF photometry of JWST/MIRI data may not be consistent with each other and recommend using similar methods as reported in previous publications. This discrepancy may decrease over the course of the JWST mission as MIRI aperture corrections and PSF models continue to be updated.

### v2

@@ 3.5 Dust Modeling Results
The Cycle 1 JWST observations were previously modeled for the optically thin and dusty sphere case in M. Shahbandeh et al. (2023). M. Shahbandeh et al. (2023) report only a total dust mass rather than separate mass components as we do here. However, we are able to reproduce all the dust properties for the dusty sphere of silicate dust case reported in M. Shahbandeh et al. (2023) by assuming since . For the optically thin case, we are unable to replicate the temperature of the hotter dust component as M. Shahbandeh et al. (2023), though we reproduce the mass of the dust (again assuming ). However, we find that our Cycle 1 value is consistent within error with our Cycle 2 value, which is well constrained by the NIR data. We note that there are no NIR spectra of SN 2017eaw at this epoch and therefore we cannot separate the SN’s NIR line emission from the hot dust emission. Given this, the temperature of the hot dust component in our Cycle 2 models should be considered an upper limit.
@@ 3.6 Dust Geometry
First, to determine the robustness of any modeled dust radii measurements, we apply the same check as S. Zs´ıros et al. (2024) and calculate the optical depth as follows (L. B. Lucy et al., 1989):
@@ 3.7 Spectral Evolution
Recent work by L. Dessart et al. (2025) presents models of a SN IIP at 1000 days post-explosion that is CSM-interacting and also contains a small mass of dust in the cold dense shell. They find there is no significant dust attenuation effect for dust interior to the ejecta due to the small angle of the inner dust relative to the emitting region of the outer ejecta. Further, in contrast to non-interacting SNe, in the L. Dessart et al. (2025) CSM-interacting SN model interior dust has no impact on the line strengths, or the hydrogen-oxygen line ratios, due to the fact that 99.7% of the model emission is from the outer ejecta. Indeed, K. E. Weil et al. (2020) note no significant signs of dust attenuation in SN 2017eaw at 900 days post-explosion. At 2668 days post-explosion, the velocity of the ejecta (7000 km/s) is similar to that of the 1000 day model (8000 km/s) in L. Dessart et al. (2025) and it is possible that the observed line attenuation is not due to interior dust. Although this physical picture may still be valid, we caution against direct comparisons given that both the inner and outer regions have evolved for an additional 1700 days. The velocity of the ejecta at 2668 days suggests that the outer regions of the ejecta have somewhat slowed due to CSM interaction and the conclusions from 1000 days may no longer be valid. L. Dessart et al. (2025) also investigate the effect of , , and of dust in the ejecta at a velocity coordinate of 8000 km/s. These models produce H profiles very similar to the one at 2668 days. The model profiles include a dip in the center (near rest wavelengths) due to the increased optical depth of the limbs of the shell. We refrain from definitively linking a similar feature in the observed H profile with dust given the low SNR of the spectrum. Further modeling of the effects of dust in CSM-interacting SNe II at late times (1000 days post-explosion) is needed to understand the evolution of H and the strongest oxygen lines.
@@ 4.1 Collisionally Heated Pre-existing Dust
The mid-infrared SED of SN 2017eaw is somewhat similar in shape to that of SNe 1987A and 1980K at significantly later phases, see Figure 10. In the case of both SNe 1987A and 1980K, the 8-20 m emission is consistent with 150-180K silicate dust, remarkably similar to SN 2017eaw. Notably the mid-infrared flux in SN 2017eaw, which is primarily dominated by the cool dust component, is almost identical in shape to Spitzer IRS spectra of SN 1987A between 6000 and 8000 days post-explosion. This component in SN 1987A has been linked to the collisional heating of the equatorial ring (E. Dwek et al., 2010; R. G. Arendt et al., 2016, 2020), and a similar scenario was suggested for SN 1980K (S. Zs´ıros et al., 2024). The mass of dust in the 160K component of SN 2017eaw is 1 order of magnitude larger than observed in SN 1987A and 1 order of magnitude smaller than observed in SN 1980K.
To determine if it is plausible for the cooler component of SN 2017eaw’s mid-infrared dust to be collisionally heated via interaction between the ejecta and CSM, we follow the method of O. D. Fox et al. (2010) for estimating the mass of dust processed by the forward shock (see also O. D. Fox et al., 2011; S. Tinyanont et al., 2016; S. Zs´ıros et al., 2022, 2024). Any pre-existing dust must reside outside the evaporation radius, inside which the peak luminosity of the SN will have destroyed any pre-existing dust grains. Assuming the temperature and peak bolometric luminosity measured by T. Szalai et al. (2019b), 14,000K and 1043 erg/s (rounded from L erg/s) respectively, the evaporation radius (Revap) is R⊙. Revap is significantly less than the ejecta radius at 2,000 days. In both JWST epochs, the SN ejecta has far surpassed the evaporation radius and could feasibly be interacting with CSM containing pre-existing dust.
@@
where is the shock velocity, is the time since explosion, and is the dust grain size. Similar to S. Zs´ıros et al. (2024), we use 5,000 km s-1 and 15,000 km s-1 and 0.005 and 0.1 m (we assume m in our dust modeling) as our lower and upper bounds, respectively. Assuming these values, the range of dust masses that could be collisionally heated is . The total dust masses observed in both Cycle 1 and 2 are in the middle of this range. We note that the velocity of the ejecta measured from the nebular spectra, 7000 km s-1, with a 0.1 m dust grain could result in a collisionally heated dust mass of . This result suggests that the majority of the dust observed in the mid-infrared could reasonably be collisionally heated pre-existing dust.
@@ 4.2 Radiatively Heated Pre-existing Dust
A simple infrared echo model, assuming the light from the SN explosion excites pre-existing dust, places the echo radius at , where is the duration of the light echo (M. F. Bode & A. Evans, 1980; E. Dwek, 1983). Assuming a lower limit of days in order for the echo to still be detectable in the Cycle 2 mid-infrared observations, this gives , significantly above the outer dust radii given by the non-optically thin models in Section 3.4. When the ejecta is between the evaporation and the light echo radii, as is the case for SN 2017eaw, the luminosity from CSM-ejecta interaction heats the pre-existing grains creating a CSM echo (O. D. Fox et al., 2010, 2011)777Sometimes referred to as a circumstellar shock echo. M. Shahbandeh et al. (2023) found that for SN 2017eaw the optical luminosity necessary to heat the grains to the temperature of the cool dust component in the Cycle 1 data exceeds the observed optical luminosity. Unsurprisingly, we find this to be the case for the Cycle 2 dust as well.
@@ 4.4 Implications for Dust Formation in CCSNe
If the mid-infrared dust emission in SN 2017eaw is primarily due to pre-existing dust, then its location on the dust formation timeline may be significantly different than shown in Figure 12. It is possible that many of the early time mid-infrared dust measurements of CCSNe are similarly contaminated with pre-existing dust. Late time dust mass measurements of SN 1987A were done in the far-IR and sub-mm and probed dust significantly colder than can be observed with JWST (M. Matsuura et al., 2011; R. Indebetouw et al., 2014; M. Matsuura et al., 2015). For SN 1980K, reported dust measurements were measured by modeling the dust attenuation on optical spectral lines. The mid-infrared SED of SN 1980K reveals 100 times less dust than indicated by the line profiles (S. Zs´ıros et al., 2024). This suggests that the majority of the dust in SN 1980K is also too cold to be observed by JWST/MIRI. The same might be true for SN 2017eaw but the signal to noise of the recent spectra makes modeling of the line profiles difficult.
@@ 6 Acknowledgments
The complete set of HST and JWST data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/tk94-1430 (catalog doi: 10.17909/tk94-1430).
This work is based in part on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from MAST at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program GO3295 and GO2666.
@@
In order to measure the flux in the MIRI images, we attempted several different photometric techniques which we compared to the published Cycle 1 photometry in M. Shahbandeh et al. (2023). In this work, we report JWST MIRI PSF photometry of SN 2017eaw as this methodology resulted in values that are most consistent with previously published photometry. M. Shahbandeh et al. (2023) did PSF photometry on the stage 2 products, a method similar to the one we use for our photometry in Section 2.1. As shown in Figure 13 (purple diamonds and pink squares), the most significant offsets between our space_phot PSF photometry and the published photometry are at 18 and 21 µm, both of which are filters where the photometric calibration maps were significantly updated888This work uses version 0056, details can be found at https://jwst-crds.stsci.edu/browse/jwst_miri_photom_0056.rmap between the publication of M. Shahbandeh et al. (2023) and the completion of this work, therefore this offset is unsurprising.
@@
Due to the discrepancy between the aperture and PSF photometry regardless of methodology, and the existence of published JWST/MIRI PSF photometry of SN 2017eaw, we opt to report only the PSF photometry discussed in Section 2.1 in this work. We note that M. Shahbandeh et al. (2023) subtract a background image, made by taking the sigma-clipped average of the individual dithers, from the level 2 images before doing PSF photometry. We opt to not do this since space_phot fits the background alongside the PSF model and generally our photometry is consistent with M. Shahbandeh et al. (2023). Importantly, we do find that, regardless of the photometric method, the total mid-infrared flux of SN 2017eaw has decreased from Cycle 1 to Cycle 2. We caution that the aperture and PSF photometry of JWST/MIRI data may not be consistent with each other and recommend using similar methods as reported in previous publications. This discrepancy may decrease over the course of the JWST mission as MIRI aperture corrections and PSF models continue to be updated.